OpenStudy (anonymous):

To get to work, Truman waited 18 minutes for the bus, then traveled 13.44 miles on the bus at a speed of 28 miles per hour. How much time passed from the time Truman left for work to the time he arrived at work? 42 minutes 28.8 minutes 18.48 minutes 46.8 minutes

OpenStudy (texaschic101):

Time = distance / speed time = 13.44/28 time = 0.48 then add your 18 minutes for the time he waited for the bus.. 18 + 0.48 = 18.48 minutes

OpenStudy (anonymous):

It was D but thanks anyway!! :)

OpenStudy (texaschic101):

I am so sorry....how did they get D ?